Circular No. 9073

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 R2 (PIGOTT-LINEAR-KOWALSKI) Further to IAUC 9072, elements by B. G. Marsden for epoch 2003: Epoch = 2003 Feb. 10.0 TT

T = 2003 Feb. 1.2551 TT  ω = 357.^{o}0844
e = 0.480777  Ω = 55.1940 }2000.0
q = 1.915714 AU  i = 46.2617

a = 3.689579 AU n^{o} = 0.1390717 P = 7.09 years On IAUC 8044, the suggestion was made that P/2003 A1 = 1783 W1,

although it was not possible unequivocally to demonstrate this at the time. Running the above orbit back to 1783 (closest approach to Jupiter 0.19 AU in 1971) yields

Epoch = 1783 Nov. 10.0 TT

T = 1783 Nov. 23.1803 TT  ω = 354.^{o}0531
e = 0.578766  Ω = 58.5969 }2000.0
q = 1.458960 AU  i = 45.7727
a = 3.463539 AU n^{o} = 0.1529058 P = 6.45 years

This result compares very favorably with the orbit by C. H. F. Peters (1860, Brünnow’s Astron. Notes, No. 19), notably with T, q, and Ω agreeing to 2.275 days, 0.0003 AU, and 0^{o}.08, respectively. Although the quality of the observations is poor, the above orbit by Marsden satisfies fourteen of them made during 1783 Nov. 22–Dec. 4 to within < 10′ in each coordinate, provided that T is corrected to 1783 Nov. 19.59.

COMET 88P/HOWELL J. Crovisier, P. Colom, N. Biver, and D. Bockelée-Morvan, LESIA,

Observatoire de Paris, report that observations of the 18-cm lines of OH in comet 88P/Howell with the Nançay radio telescope show an increase in the average OH-production rate from 1.0 ± 0.2 × 10^{28} molecules/s during Aug. 1–13 to 4.3 ± 0.5 × 10^{28} molecules/s during Aug. 28–Sept. 3.

COMETS 223P/SKIFF, 224P/LINEAR-NEAT, 225P/LINEAR Comet P/2009 L18 = 2002 S1 (cf. IAUC 9066) has been assigned the

permanent number 223P. Comet P/2009 Q2 = 2003 XD_{10} (cf. IAUC 9068) has been assigned the permanent number 224P. Comet P/2009 Q3 = 2002 T1 (cf. IAUC 9068) has been assigned the permanent number 225P. (See also MPC 66864).

2009 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green